I wanna know what the size of the largest square is that could fit in a circle with a diameter of 1".
Answer by Alan3354(69443)

I wanna know what the size of the largest square is that could fit in a circle with a diameter of 1".
---------------
The diagonal of the square is the diameter = 1.
d^2 = s^2 + s^2 = 1
2s^2 = 1
s^2 = 1/2
s = sqrt(1/2) = sqrt(2)/2 = ~ 0.70711" (length of sides)